### v1

@@ 3.3 Initialisation
As it is standard in competing approaches i.e. [40], we can use single view PS (at each view) in order to get normal and depth estimates. We start by computing per view normal maps using the state-of-the-art PS normal estimation network [9] and also use numerical integration [35] in order to get approximate depth maps. The depth maps do need to be accurate as they are only used to initialise the ray search spaces. The normal maps can be used in order to provided an additional training signal.
The SDF newtwork is initialised with weights that approximate the SDF of a perfect sphere. to speed up convergence, we always run 3 epochs with normal and silhouette loss without rendering (which is a lot more computationally expensive).
@@ 3.4 Losses
Rendering loss. We use L1 loss on the rendered intensities, i.e. $L_{rend}=|i_{rend}-i_{gt}|$. We note that to better balance the rendering data, each light source is scaled so as the maximum GT intenisty is 1 (saturated). This is performed because some of the DiliGenT-MV lights are very dark. The relative weighting of this loss when used is 1e3.
@@
Silhouette loss. As it is standard in volumetric SDF approaches, we apply binary cross entropy loss to the predicted silhuettes to the ground truth ones. We generate predicted silhuettes with the total accumulated opacity $\alpha_{total}=\prod_{j}(1-\alpha_{j})$ (which should be 1 for foreground and 0 for background points), $L_{sil}=\text{BCE}(\alpha_{total},mask)$. We consider a 5 pixel band outside the provided segmentation masks for background points. We note that we ignore the utmost pixels of the provided segmentation masks and also the innermost pixels of the background (leaving 2 pixel boundaries between foreground and background). This is done for better numerical stability as the exact discritisation of the boundaries is not very reliable. The relative weighting of this loss is 1e1.
@@
Normal loss. We apply angular normal loss to match SDF computed normals $\boldsymbol{\mathbf{N}}_{s}$ to the network predicted normals $\boldsymbol{\mathbf{N}}_{n}$. We follow [25, 26] and use angular error (as opposed to just L2 used in approaches like [40]) $L_{n}=|\text{atan2}(||\boldsymbol{\mathbf{N}}_{n}\times\boldsymbol{\mathbf{N}}_{s}||,\boldsymbol{\mathbf{N}}_{n}\cdot\boldsymbol{\mathbf{N}}_{s})|$ For rays where the accumulated opacity is less than 0.01 (i.e the ray does not intersect any surface), no normal loss is applied. In addition, following previous works, the normal loss is weighted by the obliqueness of each point $(\boldsymbol{\mathbf{N}}\cdot\boldsymbol{\mathbf{V}})$ and that stops the optimisation to try to fit occlusion boundaries which are numerically unstable. The relative weighting of this loss when used is 1.
@@
Eikonal loss. To enforce the Eikonal equation for all ray samples, we apply L1 loss which is the standard in most SDF approaches i.e. $L_{eik}=|~{}||\nabla d||-1|$. The relative weighting of this loss is 1e1.
Free space regulariser. It is important to use some sort of a regulariser to minimise free space artifacts on the inside and outside of the object, in unseen regions. To do that, at each ray, the SDF needs to be encouraged to get more negative on the points behind the first ray/surface intersection, with penalty weight the different of the target sign $s_{t}$ (-1 for inside, +1 for outside) to the soft sign i.e. $\text{softsign}(d)=\frac{d}{1+|d|}$ So we compute $L_{reg}=|\text{softsign}(d)-s_{t}|$. The is graphically shown in Figure 3 where $\boldsymbol{\mathbf{X}}_{r1}$ is infront of $\boldsymbol{\mathbf{X}}_{A}$ and thus has $s_{t}=1$, $\boldsymbol{\mathbf{X}}_{r6}$ is behind and thus has $s_{t}=-1$. Finally, to minimise the effect of regulariser on the surface evolution, we only apply it to points where their opacity contribution is less than 1% (i.e. far away from $\boldsymbol{\mathbf{X}}_{A}$, colored red on Figure 3). The relative weighting of this loss is 1e-1.
@@ 4.1 Datasets
DiLiGenT-MV. Our main evaluation target is DiLiGenT-MV [21] containing 5 objects with 96 lights in 20 views. Images are $612\times 512$ resolution with objects actually occupying a maximum of $400\times 400\text{px}$. GT meshes, camera instrisics and extrinsics and normal maps are provided, as well as point light positions and far-field light brightness ($\phi$). We note that these brightness were measure from the intensity of a flat calibration target roughly positioned at the object, so intrinsic brightness is recovered by multiplying with inverse distance square. We note that as these calibration data are unavailable, the $\phi$ is expected to be fairly inaccurate and thus is also optimised during training.
Dataset anomalies. In order to have the most fair assesement, we report the following dataset anomalies on DiLiGenT-MV and our attempts to overcome them. Firstly, the provided GT normal maps are incompatible with the GT meshes as well as the provided intrinsics and extrinsics. The best compatibility was achieved by using the intrinsic matrix for Reading for all objects while simultaneously performing SDV to fix the non-one determinant on the provided rotation matrices (especially for Reading). Despite that, there is still impossible to get a perfect match between provided normals and meshes so the exact parameters remain uncertain. For all experiments expect the first line of Table 2, we use the modified parameters (Reading instrisics and endorsing unit determinant on rotation matrices R with SVD).
In addition, the provided segmentation masks in Bear and Cow contained holes that had to be closed manually otherwise the silhouette loss would make big holes on the reconstructed meshes.
Finally, Ikehata in [11] first noticed that the first 20 images of the Bear appear to be corrupted. We also found more similarly corrupted images on other views (more visualisation in the supplementary) and did out best effort to manually mark and ignore them but it is possible that more image corruptions are still unnoticed.
Synthetic DiLiGenT-MV. To better demonstrate the effectiveness of our method without the real data corruptions discussed above, we rendered a synthetic version of DiLiGenT-MV with Blender. See Figure 4. We use the exact same objects, with the exact same poses and rendered the 96 points lighst. The objects materials where chosen to loosely mimic real objects and the albedo was set to a random pattern. Finally, we note that this synthetic data can be used to visualise shadow and indirect reflection maps which are really hard to correctly evaluate on real data. This is explored in the supplementary.
@@ 4.2 Hyperparameters and Training
We use a tensorflow port of SIREN for all the experiments. The SDF network is set to 5x512 layers (1.05M parameters) and the albedo and self reflection network to 3x256 (133K parameters). The BRDF network is set to 3x32 layers (2.5K parameters). We use 64 ray samples in a 100mm ray range and an additional 64 around the average intersection (in a shrinking distance range up to 10mm) and 64 extra samples computed with one step of Newton method (for approximating the 0 of the SDF). For each shadow ray we used 16 samples. We train with batch size of 512 rays for 100 epochs which takes approx. 20h on a NVIDIA TITAN RTX and 17GB of RAM when rendering all 96 lights. Note that the 6 lights versions completes in only 5h.
@@ 4.3 Evaluation protocol
We evaluate our method by computed Chamfer distance (marked as surface error SE) of the reconstructions and the ground truth . This is computed as the average of Hausdorf distance from reconstruction to ground truth and the opposite, with the distances computed with with Meshlab. We note that in order to have a fair comparison and not bias the error with unseen bottom of the objects, the bottom 6mm of GTs and all reconstructions are removed. All DiLiGenT-MV objects are aligned to be touching the XY plane so the cropping is straightforward. This cropping also avoids large error at some parts of the bottom of the objects that are occluded by the background (e.g. the feet of Reading, see supplementary).
Competing approaches. We compare against DiLiGeNT [21], PS-NeRF [40], MVAS [2] and SuperNormal [1]. For all of the methods, the output reconstructions were downloaded from the SuperNormal project page [^1] https://xucao-42.github.io/homepage/. Synthetic data is only used to ablate our own method as the DiLiGeNT and SuperNormal code is not available (and our method using only norma loss is very similar to PS-NeRF and SuperNormal). We also note that PS-NeRF and MVAS are also using a modified/pre-procesed version of DiLiGeNT and this is potentially alleviating some of the anomalies (especially the segmentation mask holes).
@@ 5 Experiments
Figure 4: Qualitative visualisation of re-rendering and rendering errors for synthetic and real data. We note that the scaling of the error map sets red to 0.1.
@@ 5.1 Synthetic data
As mentioned in Section 4.1 the original DiLiGenT-MV dataset contains several anomalies making hard to correctly ablate the several steps of our method. Thus ablation experiments are performed on Synthetic-DiLiGenT-MV dataset described in Section 4.
Method Bear Buddha Cow Pot2 Reading Average SE
GT Normals 0.06 0.08 0.03 0.04 0.02 0.05
Ours - [normals. only] 0.17 0.14 0.10 0.14 0.16 0.14
Ours - [intensities only] 0.13 0.29 0.07 0.20 0.11 0.16
Ours - [normals. + intensities.] 0.11 0.17 0.04 0.18 0.11 0.12
Table 1: Ablation study on DiLiGenT synthetic data. Using normals and intensities for shape estimation achieves the best error. Particularly note that it outperforms both other configurations on Bear and Cow objects indicating the the combined approach is better than a simple interpolation between the two.
We first show that our network can achieve a very low error of 0.05mm when using the ground truth normals which is not the case for the real DiLiGenT-MV dataset as shown in Table 2. We also show that if predicted normals are used the performance is worse but still quite good (ie. 0.14mm). Using intensities only achieves overall accuracy of 0.16mm. In addition, using intensities only significantly outperforms the normals only network on Bear, Cow, Reading but it is inferior on Buddha and Pot2. The reason for such a differing performance is the distribution of self reflection that affect the rendering performance more than the normal estimation network. The combined normals and intensities experiment achieves the best accuracy of 0.12mm average error. In addition, on Bear and Cow objects, the accuracy using both losses is better than any of the individual ones (0.17, 0.13 vs 0.11 for Bear) and (0.10, 0.07 vs 0.04 for Cow) therefore showing the ability to combine best of both training signals.
5.2 Real data
In this section we report our results on DiLiGenT-MV benchmark. We first show that there some issues with DiLiGenT-MV intrinsic camera parameters and/or ground truth normals. In particular, we compute the shape estimation error by using original intrinsic parameters and obtain a significantly higher error of estimated shape than on the synthetic DiLiGenT-MV benchmark. However if we take the intrinsic parameters from Reading object and use it for all objects, the error is drastically reduced (0.11 to 0.03). We note that for the Reading object, the difference of 0.06 to 0.05 is explained by adjusting the rotation matrices with SVD.
When looking at our network performance we show that we achieve state of the art results, only matching to the original DiLiGenT-MV benchmark and significantly outperform deep learning based competitors (ie. PS-NeRF [40] and MVAS [2]). Note in the original paper of PS-NeRF the performance was compared by including the bottom (invisible) part of the object which gave misleading results of deep learning method outperforming the classical one. We believe that the classical method performs so well as the objects have relatively simple geometry (especially the Cow where is mostly convex and has the least amount of shadows and self reflection). The reconstructed meshes are shown in Figure 5.
Figure 5: Qualitative results on real DiLiGenT-MV benchmark. The error bars are set to red corresponding to 1mm. We note that we are achieving consistent, uniform accuracy on all regions of all objects, including the concavity in the middle of Reading.
Note we also show that if we reduce the number of lights available (e.g. 6) and hence decrease the quality of normals reduces and the impact of rendering increases, justifying our key contribution. If a more challenging multi-view photometric stereo benchmark were to be available our approach would have display a larger gain.
Method Bear Buddha Cow Pot2 Reading Average SE
GT normals 0.13 0.15 0.10 0.12 0.06 0.11
GT normals (corr. params) 0.02 0.06 0.01 0.02 0.05 0.03
DiLiGeNT [21] 0.22 0.33 0.08 0.21 0.25 0.22
@@ 5.2 Real data
SuperNormal [1] 0.21 0.21 0.19 0.15 0.21 0.19
Ours - [normals only] - 6 lights 0.59 0.46 0.34 0.63 0.45 0.49
Ours - [intensities only] - 6 lights 0.22 0.29 0.22 0.32 0.32 0.27
Ours - [normals + intensities] - 6 lights 0.25 0.38 0.24 0.52 0.33 0.34
Ours - [normals only] 0.29 0.19 0.17 0.19 0.29 0.22
Ours - [intensities only] 0.25 0.24 0.18 0.34 0.26 0.26
Ours - [normals + intensities] 0.21 0.19 0.17 0.20 0.22 0.20
Table 2: Results on DiLiGenT-MV [21] benchmark. For all objects we report the mean shape error as well as average shape error and average median shape error on all objects. Normals are computed using the universal PS method of [9].
Finally, we achieve slightly inferior performance to SuperNormal [1], mainly because of the Pot2 object (0.2 vs 0.15mm). However given the discovered issues with the DiLiGenT-MV benchmark and code (or the normal maps used for input) not being available for this method we refrain from drawing on concusions for the reasons why it slightly outperforms our method. One possible explanation is there are more locally corrupted images (like the specular highlights on Bear) that are interfering with are renderer but do not pose issue to their particular normal estimation network. Note that SuperNormal method itself is functionally very similar to ours when no intensities (ie SuperNormal makes use of an SDF trained with normals, silhouette and Eikonal losses) are used.

### v2

@@ 3.3 Initialisation
As it is standard in MVPS approaches, e.g. [44], we can use single view PS (at each view) in order to obtain normal and depth estimates. We start by computing per view normal maps using the state-of-the-art PS normal estimation network [12] and also use numerical integration [39] in order to get approximate depth maps. The depth maps do need to be accurate as they are only used to initialise the ray search spaces. The normal maps are used in order to provided an additional training signal.
The SDF network is initialised with weights that approximate the SDF of a perfect sphere. To speed up convergence, we always run 3 epochs with normal and silhouette loss without rendering.
@@
Figure 4: Qualitative visualisation of re-rendering and rendering errors for synthetic and real data. The scaling of the error map sets red to 0.1. We note that most of the error is concentrated on the middle of concavities as self reflection is not modeled.
@@ 3.4 Losses
Rendering loss. We use L1 loss on the rendered intensities, i.e. $L_{rend}=|i_{rend}-i_{gt}|$. We note that to better balance the rendering data, each light source is scaled so as the maximum GT intensity is 1 (saturated). This is performed because some of the DiliGenT-MV lights are very dim. The relative weighting of this loss when used is 1000.
@@
Silhouette loss. We also apply binary cross entropy loss to the predicted silhuettes to the ground truth ones. We generate predicted silhuettes with the total accumulated opacity $\alpha_{total}=\prod_{j}(1-\alpha_{j})$ (which should be 1 for foreground and 0 for background points), $L_{sil}=\text{BCE}(\alpha_{total},mask)$. We consider a 10 pixel band outside the provided segmentation masks for background points. The relative weighting of this loss is 10.
@@
Normal loss. We apply angular normal loss to match SDF computed normals $\boldsymbol{\mathbf{N}}_{s}$ to the network predicted normals $\boldsymbol{\mathbf{N}}_{n}$. We follow [29, 30] and use angular error (as opposed to L2 loss used in PS-NeRF [44]) $L_{n}=|\text{atan2}(||\boldsymbol{\mathbf{N}}_{n}\times\boldsymbol{\mathbf{N}}_{s}||,\boldsymbol{\mathbf{N}}_{n}\cdot\boldsymbol{\mathbf{N}}_{s})|$. For rays where the accumulated opacity is less than 0.01 (i.e the ray does not intersect any surface), no normal loss is applied. In addition, following previous works, the normal loss is weighted by the obliqueness of each point $(\boldsymbol{\mathbf{N}}\cdot\boldsymbol{\mathbf{V}})$ and that stops the optimisation to try to fit occlusion boundaries which are numerically unstable. The relative weighting of this loss when used is 1.
@@
Eikonal loss. To enforce the Eikonal equation for all ray samples, we apply L1 loss which is the standard in most SDF approaches i.e. $L_{eik}=|~{}||\nabla d||-1|$. The relative weighting of this loss is 10.
Curvature regulariser. To minimise floater artefacts (especially on the inside of objects) and encourage the optimisation to recover the minimum surface, it important to include some curvature regulariser encourage smoothness on the volumetric normals (i.e. SDF gradients). Computing exact analytic curvatures (via auto-differentiation) has a high computational cost an is not really required, as the objective is to only use them as a regulariser. Instead, drawing inspiration from SuperNormal [4], we use the ray samples $\boldsymbol{\mathbf{X}}_{i}$ and compute finite differences along the ray as: $curv(\boldsymbol{\mathbf{X}}_{i})\approx\frac{||\boldsymbol{\mathbf{N}}_{i}-\boldsymbol{\mathbf{N}}_{i+1}||}{||\boldsymbol{\mathbf{X}}_{i}-\boldsymbol{\mathbf{X}}_{i+1}||}$. We note that an exact curvature would require finite differences along all 3-axis, but for regularisation purposes this definition is adequate, and comes with no additional SIREN queries. Note that NeILF++ [48] also includes a similar regulariser. The relative weighting of this loss is 1.
@@ 4 Experiment Setup
Method Bear Buddha Cow Pot2 Reading Avg. SE
GT Normals 0.06 0.08 0.03 0.04 0.02 0.05
N 0.17 0.14 0.10 0.14 0.16 0.14
I 0.13 0.29 0.07 0.20 0.11 0.16
N + I 0.11 0.17 0.04 0.18 0.11 0.12
Table 1: Ablation study of our method on a synthetic replica of DiLiGenT-MV [24] benchmark. We first compute our method performance using ground truth normals in order to highlight potential issues with real DiLiGenT-MV [24] benchmark (first two rows of Table 2) where recovered shape is significantly less accurate both for our and Supernormal [4] methods (0.05mm vs 0.11mm). We also include comparison of the 3 versions of our method named N, I and N+I namely normal loss only (no rendering), rendering loss only (no normals) and combined.The combined approach achieves the best error and particularly note that it outperforms both other configurations on Bear and Cow objects indicating the the combined approach is better than a simple interpolation between the two.
@@ 4.1 Datasets
DiLiGenT-MV. Our main evaluation target is DiLiGenT-MV [24] containing 5 objects with 96 lights in 20 views. Images are $612\times 512$ resolution with objects actually occupying a maximum of $400\times 400\text{px}$. Ground truth meshes, camera intrinsics, extrinsics and normal maps are provided, as well as point light positions and far-field light brightnesses ($\phi$). We note that these brightnesses were measured from the intensity of a flat calibration target roughly positioned at the location of the imaged objects, so intrinsic brightness is recovered by multiplying with inverse distance square. We note that as such calibration data is unavailable, the $\phi$ is expected to be fairly inaccurate and thus it is optimised during training.
Dataset anomalies. In order to have the most fair assessment, we report the following dataset anomalies on DiLiGenT-MV and our attempts to overcome them.
Firstly, the provided GT normal maps and masks are incompatible with renderings of ground truth meshes when provided intrinsic and extrinsic parameters are used. Note, some of the provided rotation matrices are not orthonormal and have non-unitary determinant. We follow the approach of RNB  [2] and Supernormal [4] by first performing an explicit projection matrix computation ($P=K[R|T]$) followed by QR decomposition (using OpenCV) to obtain orthonormal rotation matrix.
Secondly, the provided segmentation masks in Bear and Cow contain holes that need to be closed manually in order to prevent the silhouette loss from introducing large holes in the reconstructed meshes.
Finally, Ikehata et al [14] first noticed that the first 20 images of the Bear appear to be corrupted. We also found more similarly corrupted images on other views (more visualisations in the supplementary) and did our best effort to manually mark and ignore them however it is possible that more image corruptions are still unnoticed.
Synthetic DiLiGenT-MV. To better demonstrate the effectiveness of our method without the real data corruptions discussed above, we rendered a synthetic version of DiLiGenT-MV with Blender. See Figure 4. We use the exact same objects, with the exact same poses and rendered the 96 points lighst. The objects materials where chosen to loosely mimic real objects and the albedo was set to a random pattern. Finally, we note that this synthetic data can be used to visualise shadow and indirect reflection maps which are really hard to correctly evaluate on real data.
@@ 4.2 Hyperparameters and Training
We use a tensorflow port of SIREN for all the experiments. The SDF network is set to 5x512 layers (1.05M parameters) and the albedo network to 3x256 (133K parameters). The BRDF network is set to 3x32 layers (2.5K parameters). We use 64 ray samples in a 100mm ray range and an additional 64 around the average intersection (in a shrinking distance range up to 10mm) and 64 extra samples computed with one step of Newton method (for approximating the 0 of the SDF). For each shadow ray we used 16 samples. We train with batch size of 512 rays for 100 epochs which takes approx. 20h on a NVIDIA TITAN RTX and 17GB of RAM when rendering all 96 lights. Note that the 6 lights 6 views version completes in only 2h.
@@ 4.3 Evaluation protocol
We evaluate our method by computed Chamfer distance (marked as surface error SE) of the reconstructions and the ground truth . This is computed as the average of asymmetric Hausdorf distance from reconstruction to ground truth and the opposite, with the distances computed with with Meshlab. We note that in order to have a fair comparison and not bias the error with unseen bottom of the objects, the bottom 6mm of GTs and all reconstructions are removed. All DiLiGenT-MV objects are aligned to be touching the XY plane so the cropping is straightforward. This cropping also avoids large error at some parts of the bottom of the objects that are occluded by the background (e.g. the feet of Reading, see supplementary). Thus our reported error (in Table 2) is generally lower than the numbers originally reported in other works.
Competing approaches. We compare against DiLiGenT-MV [24], PS-NeRF [44], MVAS [5] Brahimi et. al, [1] RNB  [2] and Supernormal [4]. DiLiGenT-MV [24] and Brahimi et. al, [1] are closed source ones so we use the meshes computed by the original authors; for all other methods we use the meshes we were able to reproduce from their codes. We note that Supernormal [4] offers the best perfomance and by far the least computational cost so it is used for ablating the sparse lights and views scenario as well. Note that Brahimi et. al, [1] was only computed on the sparse scenario therefore that is the only available comparison for them.
@@ 5 Experiments
Figure 5: Qualitative results on real DiLiGenT-MV [24] benchmark. For each mesh vertex, the minimum distance to the GT mesh is shown with the error bars set to red corresponding to 1mm. We note that we are achieving consistent, uniform accuracy on all regions of all objects, including the concavity in the middle of Reading.
@@ 5.1 Synthetic data
As mentioned in Section 4.1 the original DiLiGenT-MV [24] dataset contains several anomalies making hard to correctly ablate the several steps of our method. Thus ablation experiments are performed on Synthetic-DiLiGenT-MV dataset described in Section 4.
We first show that our network can achieve a very low error of 0.05mm when using the ground truth normals which is not the case for the real DiLiGenT-MV dataset as shown in Table 2. We also show that if predicted normals are used the performance is worse: 0.14mm. Using the rendering of intensities only achieves overall accuracy of 0.16mm. In addition, using intensities only significantly outperforms the normals only network on Bear, Cow, Reading but it is inferior on Buddha and Pot2. The reason for such a differing performance is that the presence of strong self reflection effects presents a more difficult task to the rendering network than the normal estimation network. The combined normals and intensity rendering experiment achieves the best accuracy of 0.12mm average error. In addition, on Bear and Cow objects, the accuracy using both losses is better than any of the individual ones (0.17mm, 0.13mm vs 0.11mm for Bear) and (0.10mm, 0.07mm vs 0.04mm for Cow) therefore showing the ability to combine best of both training signals.
Method Bear Buddha Cow Pot2 Reading Avg.
SpN [4] GT N 0.16 0.12 0.06 0.10 0.13 0.11
Ours GT N 0.13 0.15 0.10 0.12 0.06 0.11
DiLiGenT-MV [24] 0.22 0.33 0.08 0.21 0.25 0.22
@@
SpN [4] 0.19 0.21 0.21 0.14 0.22 0.20
RNB [2] 0.25 0.21 0.31 0.18 0.27 0.24
Ours N 0.29 0.19 0.17 0.19 0.29 0.22
Ours I 0.25 0.24 0.18 0.34 0.26 0.26
Ours N + I 0.21 0.19 0.17 0.20 0.22 0.20
Table 2: Results on original dense DiLiGenT-MV [24] benchmark. For all objects we report the Chamfer distance error as well as average error on all objects. Normals are computed using the universal PS method of [12]. We evaluate 3 versions of our method named N, I and N+I namely normal loss only (N) - no rendering, rendering loss only (I) - no normals loss and combined (N+I). We also include comparison with Supernormal (SpN [4]) with ground truth normals (GT N).
5.2 Real data
In this section we report our results on DiLiGenT-MV [24] benchmark in both original dense setup (20 views, 96 lights) and various sparse setups.
Method Bear Buddha Cow Pot2 Reading Avg.
SpN [4] [6,6] 0.48 0.67 0.27 0.39 1.21 0.61
Ours N + I [6,6] 0.32 0.35 0.28 0.33 0.63 0.38
Brah. et al [1] [6, 75] 0.38 0.32 0.24 0.29 0.47 0.34
SpN [4] [6, 30] 0.29 0.25 0.17 0.17 0.75 0.33
Ours N + I [6, 30] 0.30 0.29 0.16 0.25 0.25 0.25
SpN [4] [20, 6] 0.28 0.53 0.28 0.30 0.36 0.35
Ours N + I [20, 6] 0.25 0.38 0.24 0.52 0.33 0.34
Table 3: Results on sparse DiLiGenT-MV [24] benchmark. We include three sparse cases (marked with [views, lights] and compare with SpN [4] which is the best performing dense competitor. We also include the comparison with Brahimi et al [1] which was computed with 6 views and 75 lights which is comparable to our 6 views 30 lights case. We note that we significantly outperform Supernormal [4] (0.38mm vs 0.61mm) on the most sparse setup [6 views, 6 lights] and also confidently outperform Brahimi et. al [1] (0.25mm vs 0.34mm) despite having a significantly smaller number of lights in [6 views, 30 lights] setup. Supernormal [4] matches our performance when more views are used.
Original (dense) DiLiGenT-MV setup. Our method achieves (0.2mm) state of the art results (see Table 2), only matching to the original DiLiGenT-MV benchmark (0.22mm) and Supernormal [4] (0.2mm). Note we significantly outperform most deep learning based competitors (ie. PS-NeRF [44], MVAS [5] and RNB [2]). Also note in PS-NeRF [44] the performance of was compared by including the bottom (invisible) part of the object which gave misleading results of the deep learning method outperforming the classical one proposed originally with the DiLiGenT-MV [24] benchmark. We believe that the classical method performs so well as the objects have relatively simple geometry (especially the Cow where is mostly convex and has the least amount of shadows and self reflection). The reconstructed meshes are shown in Figure 5.
Sparse DiLiGenT-MV setup. In Table 3 provide our results on various sparse setups of DiLiGenT-MV [24] benchmark. In particular, we include three sparse cases (marked with [views, lights]). In all cases we compare with Supernormal [4] which is the best competitor on dense DiLiGenT-MV benchmark and has code available. We also include the comparison with Brahimi et al [1] (code is not available) which was originally computed with 6 views and 75 lights which is comparable to our 6 views 30 lights case. We note that we significantly outperform Supernormal [4] (0.38mm vs 0.61mm) on the most sparse setup [6 views, 6 lights] where the normal estimates are very inaccurate (see Figure 1) and confidently outperform Brahimi et. al [1] (0.25mm vs 0.34mm) despite having a significantly smaller number of lights in [6 views, 30 lights] setup. Supernormal [4] matches (0.34 mm vs 0.35mm) our performance when more views (20) are used.
It is also noteworthy that only using 6 lights and 6 views (around 2% of the total data) only increases the total error from 0.2mm to 0.38mm signifying the need for a more challenging multi-view photometric stereo benchmark where the extra images would be required to achieve sub-millimeter accuracy. Additionally, a more challenging dataset would allow newer approaches to display a larger gain vs the original method, which is still marginally worse than SOTA (0.22mm vs 0.2mm).